OpenStudy (anonymous):

The pool at "Splash Around" is open for 14 weeks during the summer. You can swim for $6 a session, or you can buy a membership for $100 and pay only $4 a session. Which equation would you use to determine how many sessions you must use the pool to justify buying the membership?

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) You want to create a system of equations. First create an equation for the first one. Then the second problem. We know it is open 14 weeks You can pay $6 a session. The equation simplifies to: 6(14) = Final cost The other equation, you're given $100 membership, and $4 a session That simplifies to 100 + 4x = final cost